how much sucrose (MW 342g/mol) you need to make 500ml of a 0.2ml of a 0.2 solution
1
Expert's answer
2019-01-29T06:36:45-0500

c= n/V

n= c*V = 0.2 mol/L * 0.5 L = 0.1 mol

n(sucrose) = 0.1 mol

m= M*n

m(sucrose) = 342g/mol * 0.1 mol = 34.2 g

Answer: 34.2 g